OpenStudy (anonymous):

PLEASE HELP! (x^2-2x-15) / (x+3) = ?

OpenStudy (anonymous):

really means "factor and cancel" since \[\frac{x^2-2x-15}{x-3}=\frac{(x-3)(x+5)}{x-3}\]

OpenStudy (anonymous):

can you show me using long division, please? that's how my teacher taught it

OpenStudy (anonymous):

i cannot write long division here no way to do it, but i can say you do it like long division with numbers

OpenStudy (anonymous):

draw it out